OpenStudy (anonymous):

hi, can someone help to just expand this ...... (e^u + e^-u)^2

OpenStudy (anonymous):

\[(e^u + e^-u)^2\]

zepdrix (zepdrix):

\[\Large \left(e^u+e^{-u}\right)^2\quad=\quad (e^u)^2+2e^ue^{-u}+(e^{-u})^2\]

zepdrix (zepdrix):

Do you understand how to simplify each term? Try to recall your rules of exponents! :)

zepdrix (zepdrix):

Some helpful rules of exponents: \[\Large \color{orangered}{(x^a)^b\quad=\quad x^{a\cdot b}}\]\[\Large \color{royalblue}{x^ax^b\quad=\quad x^{a+b}}\]

OpenStudy (anonymous):

yes, but ive got brain freeze..... is that as far as you can go with the expanding ?

zepdrix (zepdrix):

\[\Large (e^u)^2\quad=\quad e^{2u}\]Understand the first one? :o

OpenStudy (anonymous):

you did a much better job of the equation using the editor, i dont know what i did wrong ?

zepdrix (zepdrix):

If you need to put multiple things into the exponent position, you have to use the fancy curly braces {} Example: a^{3x} -> \(\Large a^{3x}\)

OpenStudy (anonymous):

i guess , yes. but it would be far simpler if the power was a number instead of the variable u

zepdrix (zepdrix):

Do you understand how the middle term simplifies? It should shrink down very nicely.

OpenStudy (anonymous):

ahhh... thx, thats good to know.. i would not have figured out the brackets :)

zepdrix (zepdrix):

\[\Large 2e^ue^{-u}\quad=\quad 2e^{u+(-u)}\quad=\quad ?\]

OpenStudy (anonymous):

thx, its 6am here, too early for me, i'll get back to it :)

OpenStudy (anonymous):

\[\Large \color{orangered}{(x^a)^b\quad=\quad x^{a\cdot b}} \Large \color{royalblue}{x^ax^b\quad=\quad x^{a+b}} \] trap for newbies :)

zepdrix (zepdrix):

-_-